Alloy Wheels inc. is a manufacturer of premium car wheels. It is considering to invest in a new production line which will cost $10m. The finance director of the company decided that a bond is is the best way to finance the project. Based on current market conditions, the bond issue will have 7-years to maturity, with a coupon of per year. The par value of the bond will be $1.000 per bond. The bond would be priced at 115 percent of its par value. The company pays a corporation tax rate of 40 percent. 


Furthermore, the stock market return offers and the risk free rate pays 1.5. The beta of the company is 1.5. 


Required:

 A Calculate the cost of debt for the company before tax?  

 B. Calculate cost of equity. 
Answer #1

Calculation of Cost of Debt

Formula:

Kd = I ( 1 - Tax Rate ) / NP x 100

Where,

Kd = Cost of Debt

I = Interest = 1000 x 6% = 60

NP = Net Proceed or Market Price = 1000 + 115% = 1150

Tax Rate = 40% or 0.40

Kd = 60 (1 - 0.40) / 1150 x 100

Kd = 3.13%

Calculation of Cost of Equity

Formula

Ke = Rf + (Rm - Rf ) x Beta

Where,

Ke = Cost of Equity

Rf = Risk Free Rate = 1.5%

Rm = Market Rate of Return = 8%

Beta = 1.5

Ke = 1.5 + ( 8 - 1.5) x 1.5

Ke = 11.25%